# ST-GS: Vision-Based 3D Semantic Occupancy Prediction with Spatial-Temporal Gaussian Splatting

### v1

@@ I Introduction
Comprehensive 3D scene understanding is a fundamental requirement for modern autonomous driving systems. Vision-based methods have gained increasing attention due to their cost-effectiveness and scalability compared to LiDAR-based approaches [1]. However, accurately modeling complex and irregular objects in dynamic driving scenes remains challenging for mapless driving [2]. Semantic occupancy prediction provides a promising solution by jointly estimating volumetric occupancy and semantic labels of arbitrary-shaped objects in 3D space, thereby improving the reliability and safety of autonomous driving in complex environments [3].
@@ II Related Work
3D semantic occupancy prediction has gained significant attention in recent years, as it provides a more comprehensive representation of surrounding environments compared to conventional 3D detection or segmentation tasks. This capability makes it particularly vital for autonomous driving. Early studies primarily relied on LiDAR point clouds, leveraging their precise geometric information for occupancy estimation [10, 11]. However, LiDAR sensors are expensive and often degrade under adverse weather or poor lighting conditions. To overcome these limitations, vision-based 3D semantic occupancy prediction has emerged as an active area of research. Recent works have adopted voxel-based representations to model 3D occupancy [12, 13, 14], as voxels effectively capture continuous 3D structures within a defined spatial volume. Nonetheless, the inherent sparsity of real-world 3D scenes makes voxel-based methods computationally expensive. Consequently, more efficient scene representations have been explored. TPVFormer [15] employs a tri-plane representation, which enforces stronger spatial constraints compared to BEV-based approaches [6] that project the scene onto a single plane. Despite these advances, both voxel-based and BEV-based methods require a dense grid of representations to model the 3D environment. In contrast, Gaussian-based approaches encode 3D scenes more compactly, using fewer primitives while maintaining geometric fidelity. Motivated by these advantages, our work focuses on Gaussian-based semantic occupancy prediction.
@@ III-C2 View-Guided Attention
Unlike GGA uses a predefined uniform sampling paradigm consisting of a set of directions evenly distributed in 3D space, motivated by view attention [21], we design a View-Guided Attention (VGA) mechanism. As demonstrated in Fig. 3(b), VGA generates offsets along the camera viewing directions, enabling more effective spatial information aggregation across overlapping multi-view image features by leveraging cross-view geometric priors. Similarly, for each Gaussian instance, we initialize a set of sampling offsets $\mathcal{P}^{V_{L}}=\{\mathcal{P}^{V_{L}}_{i}\in\mathbb{R}^{3}\}_{i=1}^{M}$ on a 2D grid in the local view coordinate frame $V_{L}$ (y–z plane). To achieve instance-specific adaptivity, we also predict an offset for each Gaussian instance from its embedding. The local sampling offsets $\Delta{\mathcal{P}}^{V_{L}}$ can then be obtained by:
@@ III-D2 Gated Temporal Feature Fusion
where $\texttt{MLP}(\cdot)$ represents a multi-layer perception block.
@@ IV-B1 Main Results
Table I provides a comprehensive comparison of our ST-GS with existing methods on the nuScenes validation split, with the best and second-best results highlighted in bold and underlined, respectively. Our ST-GS consistently outperforms both previous voxel-based approaches and recent Gaussian-based methods. Compared to our baseline, GaussianFormer [8], ST-GS achieves notable improvements of 10.22% in IoU and 12.20% in mIoU. Even against its successor, GaussianFormer-2 [9], ST-GS delivers substantial gains of 7.59% in IoU and 7.04% in mIoU. Moreover, our method attains the highest performance across most semantic categories. These results demonstrate the effectiveness of our framework and underscore that enhancing spatial interactions and incorporating temporal information are both crucial for accurate 3D semantic occupancy prediction.

### v2

@@ I Introduction
Comprehensive 3D scene understanding is a fundamental requirement for modern autonomous driving systems. Vision-based methods have gained increasing attention due to their cost-effectiveness and scalability compared to LiDAR-based approaches [1]. However, accurately modeling complex and irregular objects in dynamic driving scenes remains challenging for mapless driving [2]. Semantic occupancy prediction provides a promising solution by jointly estimating volumetric occupancy and semantic labels of arbitrary-shaped objects in 3D space, thereby improving the reliability and safety of autonomous driving in complex environments [3, 4].
@@ II Related Work
3D semantic occupancy prediction has gained significant attention in recent years, as it provides a more comprehensive representation of surrounding environments compared to conventional 3D detection or segmentation tasks [11]. This capability makes it particularly vital for autonomous driving [12]. Early studies primarily relied on LiDAR point clouds, leveraging their precise geometric information for occupancy estimation [13, 14]. However, LiDAR sensors are expensive and often degrade under adverse weather or poor lighting conditions. To overcome these limitations, vision-based 3D semantic occupancy prediction has emerged as an active area of research. Recent works have adopted voxel-based representations to model 3D occupancy [15, 16, 17, 18], as voxels effectively capture continuous 3D structures within a defined spatial volume. Nonetheless, the inherent sparsity of real-world 3D scenes makes voxel-based methods computationally expensive. Consequently, more efficient scene representations have been explored. TPVFormer [19] employs a tri-plane representation, which enforces stronger spatial constraints compared to BEV-based approaches [7] that project the scene onto a single plane. Despite these advances, both voxel-based and BEV-based methods require a dense grid of representations to model the 3D environment. In contrast, Gaussian-based approaches encode 3D scenes more compactly, using fewer primitives while maintaining geometric fidelity. Motivated by these advantages, our work focuses on Gaussian-based semantic occupancy prediction.
@@ III-C2 View-Guided Attention
Unlike GGA, which utilizes a predefined uniform sampling paradigm consisting of a set of directions evenly distributed in 3D space, motivated by view attention [25], we design a View-Guided Attention (VGA) mechanism. As demonstrated in Fig. 3(b), VGA generates offsets along the camera viewing directions, enabling more effective spatial information aggregation across overlapping multi-view image features by leveraging cross-view geometric priors. Similarly, for each Gaussian instance, we initialize a set of sampling offsets $\mathcal{P}^{V_{L}}=\{\mathcal{P}^{V_{L}}_{i}\in\mathbb{R}^{3}\}_{i=1}^{M}$ based on a 2D grid in the local view coordinate frame $V_{L}$ (y–z plane). To achieve instance-specific adaptivity, we also predict an offset for each Gaussian instance from its embedding. The local sampling offsets $\Delta{\mathcal{P}}^{V_{L}}$ can then be obtained by:
@@ III-D2 Gated Temporal Feature Fusion
where $\texttt{MLP}(\cdot)$ represents a multi-layer perceptron block.
@@ IV-B1 Main Results
Table I offers a comprehensive comparison of our ST-GS with existing methods on the nuScenes validation split, with the best and second-best results highlighted in bold and underlined, respectively. Our ST-GS consistently outperforms both previous voxel-based approaches and recent Gaussian-based methods. Compared to our baseline, GaussianFormer [9], ST-GS achieves notable improvements of 10.22% in IoU and 12.20% in mIoU. Even against its successor, GaussianFormer-2 [10], ST-GS delivers substantial gains of 7.59% in IoU and 7.04% in mIoU. Moreover, our method attains the highest performance across most semantic categories. These results demonstrate the effectiveness of our framework and underscore that enhancing spatial interactions and incorporating temporal information are both crucial for accurate 3D semantic occupancy prediction.